OpenStudy (anonymous):

5- a-1 over a+2=a^2-1 over a+2 5- (a-1/a+2)=a^2-1/a+2

OpenStudy (pottersheep):

Hmm

OpenStudy (pottersheep):

Well thhey have the same denominator

OpenStudy (anonymous):

I have gotten up to 0=a^2+4a-13, but I believe I went wrong at some point, because it should come out as a whole number..

OpenStudy (pottersheep):

Ok let me try it out

OpenStudy (pottersheep):

i got 5 - a - 1/ a^2 -1

OpenStudy (pottersheep):

WAIT are you simplifying???

OpenStudy (anonymous):

I am factoring

jhonyy9 (jhonyy9):

5-(a-1)/(a+2) = (a2 -1)/(a+2) 5(a+2)-(a-1)=a2 -1 5a +10 -a +1 =a2 -1 4a +11 =a2 -1 a2 -1 -11 -4a =0 a2 -4a -12 =0 (a-6)(a+2)=0 a_1= -2 and a_2= 6